OpenStudy (anonymous):

A radioactive source has a half life of 30 days. During a period of 90 days, the fraction of atoms that would have decayed would be what

OpenStudy (shamim):

so u wanna know the amount of atoms will b decayed. is not it

OpenStudy (shamim):

the formula is\[N=N _{o} e ^{ -\lambda t}\]

OpenStudy (shamim):

here N= number of atoms remaining after 90 days t= 90 days

OpenStudy (shamim):

\[N _{o}=\] total number of atoms

OpenStudy (shamim):

\[\lambda =\] decay constant

OpenStudy (shamim):

\[\lambda =\frac{ 0.693 }{ T _{\frac{ 1 }{ 2 }} }\]

OpenStudy (shamim):

here \[T _{\frac{ 1 }{ 2 }}=\] half life of radioactive material

OpenStudy (shamim):

now u can try to solve it urself

OpenStudy (anonymous):

Thanks Prof Shamin. I have my answer already

OpenStudy (anonymous):

A little easier approach for this problem is to realize that with a half-life of 30 days 90 days represent 3 half-lives of decay. you have 1/2 of 1/2 of 1/2 left ie 1/8 of the original amount is left meaning 1-1/8 =7/8 would have decayed

OpenStudy (anonymous):

Nice one gleem